OpenStudy (anonymous):

A line with a slope of −5 passes through the point (4, 8).

OpenStudy (campbell_st):

the slope intercept form is y = mx + b m = slope and b = y intercept so the line is y = -5x + b to find the value of b, substitute the values of x and y from the point.

Directrix (directrix):

y = -5x + b 8 = -5(4) + b 8 = -20 + b 28 = b ---------- y = -5x + 28